Given Qd = 60 – 15P + P², the (point) price elasticity of demand at a price of 5.


1
Expert's answer
2021-05-16T19:22:37-0400

Qd = 60 – 15P + P².

At a P = 5:

Qd = 60 - 15×5 + 25 = 10.

Q'd = -15 + 2P = -15 + 2×5 = -5.

The (point) price elasticity of demand is: "Ed = -5\u00d75\/10 = -2.5."

So, the demand is elastic.